# Comparison of renormalized interactions using one-dimensional few-body systems as a testbed

### v1

@@ I Introduction
Experimental advances in cold-atom physics allowed one to create and study (quasi) one-dimensional (1D) systems of a few particles in a highly controllable setting [1, 2, 3]. This rapid progress fueled studies of the corresponding Schrödinger equation [4, 5]. Thanks to the small number of particles, this equation can be solved accurately using numerical methods, such as the multi-layer multi-configuration time-dependent Hartree method for atomic mixtures [6, 7, 8], flow equation approaches [9, 10, 11], coupled cluster methods [12, 13, 14], or exact diagonalization methods, also known as the configuration interaction method (CI) [15, 16, 17, 18, 19, 20, 21, 22, 23] and its extension, the importance truncated CI [24, 25]. These powerful numerical techniques, however, usually face problems for strong interactions (see, e.g., [17, 13]).
@@
Although zero-range interactions require no regularization in one-dimensional geometries (unlike in 2D and 3D), Refs. [30, 31, 32, 33, 34, 35, 36, 37] demonstrate that renormalized interactions significantly boost the convergence of numerical simulations, enabling calculations of larger particle numbers and stronger interactions. However, to the best of our knowledge, these potentials were only benchmarked against the bare interaction. A direct comparison between different renormalization methods is still lacking and this paper aims to fill this gap. Additionally, we discuss a unitary transformation of observables together with the potential. This procedure is well-known in nuclear physics, but not in one-dimensional few-body problems. We will showcase that such a transformation can be necessary to improve the convergence of observables other than the energy.
@@
Our testbed for illustrating differences between different renormalization methods is a system of harmonically trapped fermions interacting via short range interactions. Many properties of this system are well-studied (see, e.g., [15, 38, 16, 17, 13, 39]), while at the same time the convergence in a harmonic oscillator is slow [13], making the system an ideal testbed for renormalized interactions and an optimal candidate for comparative studies. Furthermore, the corresponding two-body problem is analytically solvable [40, 41], which will facilitate construction of effective interactions in our calculations.
@@ II.1 System
where $x_{i}$ and $x_{j}$ are the coordinates of particles $i$ and $j$, and $V(x_{i},x_{j})$ is their interaction potential. The operators of the kinetic energy $T$ and the external trapping potential $W$ are given by
@@
where $N=N_{\uparrow}+N_{\downarrow}$ is the number of fermions. We assume that the spin of a particle is conserved so that $N_{\uparrow}$ and $N_{\downarrow}$ are fixed. As we are interested in low-energy scattering of atoms, we assume that $V$ is a zero-range (contact) potential [46, 47]. This implies that only fermions of different spins can interact as the Pauli exclusion principle forbids two fermions with equal spin to occupy the same spatial orbital.
@@ II.2 Numerical method: Configuration interaction method
where $H_{l}$ is the $l$th Hermite polynomial. In the CI method, we first construct $N-$particle basis states and then use them to build the Hamiltonian matrix, which is subsequently diagonalized using the Arnoldi/Lanczos algorithm [51].
@@ III Renormalized interaction
In this work, we renormalize the contact interaction in three different ways, i)-iii) mentioned above in Sec. II.1 (see also Fig. 1). We will benchmark the convergence of the different regularization methods against the bare contact interaction and then compare them with each other.
@@ III.1 Running coupling constant
where $E_{1+1}^{\mathrm{unconf.}}$ is the two-body energy of the unconfined system. For now, we discuss only $g<0$ because we want to use the energy of a bound system with $E_{1+1}^{\mathrm{unconf.}}<0$ as the renormalization condition. Note that in 1D, there is always a bound state if $g<0$ [55].
By solving Eq. (6) in momentum space, we arrive at the following equation for the energy (see the supplementary material [56] for details):
@@
This equation renormalizes the interaction in a finite Hilbert space, in the sense, that the ground state energy $E_{1+1}^{\mathrm{unconf.}}$ becomes independent of the size of the Hilbert space. Instead of the bare value $g$, the two-body energy $E_{1+1}^{\mathrm{unconf.}}$ now determines the physical properties of the system.
In the limit $\lambda\to\infty$, we derive from Eq. (8)
@@
where $C=g^{2}\braket{\Psi}{\delta(x)}{\Psi}$ is the contact parameter [57]. The functional behavior of Eq. (10) is in agreement with Refs. [58, 36]. We will see that, in general, the convergence behavior depends on the trapping potential. In an unconfined system the cutoff $\lambda$ is a continuous parameter while for the trapped system the cutoff $n$ is discrete. We connect $\lambda$ and $n$ below using the semi-classical approximation, which is limited to large quantum numbers [^3] Large means that the energy of an excited state is much larger than the excitation energy to a neighboring state of a higher energy. For the harmonic oscillator, which we discuss below, this means that $n\gg 1$. A different line of argument with the same conclusion can be based on the asymptotic approximation of Hermite polynomials, see, e.g., Ref. [59]..
Figure 2: Running coupling constant for $g(n\to\infty)=-2.5607$ (corresponding to $E_{1+1}=-1$, the two-body energy in the trapped system). Dots show $g(n)$ calculated by explicitly fixing the two-body ground state energy of two fermions in a harmonic trap for each single-body cutoff $n$. Squares show $g_{\mathrm{unconf}.}$ calculated by integrating out high momentum contributions of two fermions in free space, Eq. (8). The parameter $\lambda$ is related to the single-body cutoff $n$ via Eq. (12). The solid line represents the value of the coupling constant for an infinitely large Hilbert space, $g(n\to\infty)=-2.5607$.
@@
For a derivation of this equation and an extension to general polynomial potentials, see the supplementary material [56]. We see that convergence can be slow for a harmonic potential (see also Ref. [13]), making the use of renormalized interactions for trapped systems more important than in an unconfined case.
To compare $g(n)$ with $g_{\mathrm{unconf.}}(2\sqrt{n})$, we fix the value of $g(n\to\infty)$ and insert the expression $E_{1+1}^{\mathrm{unconf.}}=-g(n\to\infty)^{2}/4$ in Eq. (8) to calculate $g_{\mathrm{unconf.}}$. Then, we compute the two-body ground state energy of the trapped system (see Eq. (23)) and tune $g(n)$ such that this energy is reproduced in a truncated Hilbert space for each cutoff. In Fig. 2 we show the calculated running coupling constants $g(n)$ and $g_{\mathrm{unconf.}}(2\sqrt{n})$ [$g(n\to\infty)=-2.5607$ corresponding to $E_{1+1}=-1$, the two-body energy in the trapped system]. The behavior of $g(n)$ resembles that of $g_{\mathrm{unconf.}}$ for all values of $n$ qualitatively. For large values of $n$ they agree even quantitatively (less than 1% disagreement for $n\geq 21$). The disagreement for small values of $n$ can be explained by the validity of the semi-classical approximation used to derive Eq. (12), as it only holds in the limit of large quantum numbers.
@@ III.2 Effective interactions
with $T$ the kinetic energy operator. This renormalization scheme could be interpreted as a unitary transformation of the Hamiltonian [29, 60], see also supplementary material [56]. Therefore, operators should be transformed alongside the Hamiltonian transformation.
@@
Note that $\widetilde{O}^{(2)}_{\mathcal{M}\mathcal{M}}$ is represented in eigenstates of the two-body solution |$\ket{\Psi_{i}}$ while $O^{(2)}_{\mathcal{M}\mathcal{M}}$ is written in the two-body basis functions |$\ket{i}$ (cf. supplementary material [56]).
@@
Since we know the exact two-body solution of our system in terms of relative and center-of-mass coordinates (see below) the transformation of observables is straightforward if an observable can be factorized in these coordinates. However, if operators do not factorize, the transformation requires more involved calculations, for details see the supplementary material [56].
@@
with $x=(x_{1}-x_{2})/\sqrt{2}$ and $X=(x_{1}+x_{2})/\sqrt{2}$ being the relative and center-of-mass coordinates. Only the relative part of the wavefunction is influenced by the interaction and therefore $\Phi_{n_{i}}(X)$ is simply the harmonic oscillator eigenfunction; the corresponding energy is $E_{n_{i}}=n_{i}+1/2$. For the eigenfunction in relative coordinates, odd states are not influenced by the interaction and therefore the eigenfunctions are the ones of the harmonic oscillator with the same energy. For even states one derives:
@@
We use the general framework and the two-body solution presented above to build the effective interaction in two different ways. First, we construct the effective interaction in relative coordinates (i.e., using $\psi_{\nu_{i}}$ from above) and transform this interaction matrix to the laboratory frame where our calculations are performed. This is possible as only the relative part of the wavefunction is affected by the interaction. This approach is the one used in Refs. [20, 30, 31]. For the relevant equations we refer to supplementary material [56] and Ref. [20].
Note that $V_{\mathrm{eff}}^{\mathrm{rel}}$ does not lead immediately (i.e., for any cutoff parameter) to a correct full two-body spectrum in a truncated Hilbert space – the relative and center-of-mass coordinates are coupled due to the finiteness of the Hilbert space. For the low-energy part of the spectrum, this coupling is weak, and turns even weaker by increasing the cutoff parameter.
@@ III.2.2 Effective interaction from the full two-body energy spectrum: $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$
We also build the effective interaction using the full two-body solution presented in Eq. (21), see the relevant equations in supplementary material [56]. With this approach, the $1+1$ spectrum is correct for all one-body cutoffs. Note that $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ is harder to construct in comparison to $V_{\mathrm{eff}}^{\mathrm{rel}}$, in particular, for $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ the overlap between the laboratory frame and the exact two-body solution has to be calculated. However, its performance is in general better compared against our other renormalization methods as we illustrate in the next section. Furthermore, $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ is the only option for non-harmonic traps for which a factorization into relative and center-of-mass motions is not possible.
Note that effective interactions discussed above can be constructed in a similar manner for higher-dimensional systems and other (non-harmonic) traps. The prerequisite is accurate knowledge of a solution to a two-body problem, which can be obtained either analytically or numerically. A numerical solution is typically the only option for non-contact interactions and non-harmonic traps for which an analytical solution is not attainable.
@@ IV Results
Note that although we use the $1+1$ system to renormalize our interactions, not all two-body observables, for example the density, are fixed by the renormalization scheme, i.e., a large number of basis states might still be needed for convergence. Therefore, the simplest system with one spin-up and one spin-down fermion provides valuable insight into the convergence properties of the different renormalization schemes. In particular, since the $1+1$ system is exactly solvable, we can use its exact solution for accurate benchmarking. In practical calculations, the simulation of larger particle numbers is more important. We showcase the convergence of the renormalized interactions by studying a $1+3$ system with $N_{\uparrow}=1$ and $N_{\downarrow}=3$. [We observed similar convergence for other particle numbers, e.g., $2+2$.]
@@ IV.1 Energy
Figure 3: Energy of the system for different particle numbers and interactions. Panels a) and b) demonstrate the energy of the second excited state of the $1+1$ system; panels c) and d) show the ground state energy of $1+3$. Panels a) and c) are for attractive interaction with $g=-2.5067$ (which corresponds to $E_{1+1}=-1$); b) and d) illustrate the system with repulsive interactions, $g=3$ (which corresponds to $E_{1+1}=1.6$). Numerical data are shown with markers of different shape; lines are added to guide the eye. The interested reader can see these data plotted as a function of $1/n$ and also using the log-log scale for $1+1$ in the supplementary material [56].
The energies, $E(g)-E(g=0)$, obtained with the different renormalization schemes discussed above are shown in Figure 3 as a function of the one-body cutoff parameter, $n$. As a relevant scale to judge the convergence we use the harmonic oscillator energy unit, $\hbar\omega=1$; all data are shown with respect to it. Panels a) and b) present results for the second excited state [^8] We do not show the ground state energy of the $1+1$ system as this quantity is fixed for $g(n)\delta$ and $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$. of the $1+1$ system, for representative attractive and repulsive interactions, respectively. Panels c) and d) demonstrate the ground-state energy of the $1+3$ system, which does not have an analytical solution. Hence, its numerical solution is more relevant from the practical point of view.
We compare our numerical data for the $1+1$ system with the exact values, see Eq. (23). First thing to notice is that the results obtained with the bare interaction always converge slower to the limit $n\to\infty$ than those for the effective interactions. For example, for attractive interaction the energy calculated with $g\delta(x)$ has $E(n=31)-E(n=\infty)\approx 0.33$ for the highest presented cutoff; at the same time, the energy calculated with $g(n)\delta(x)$ deviates from the analytical result by $E(n=31)-E(n=\infty)\approx 0.035$. For the effective interactions we see that $V_{\mathrm{eff}}^{\mathrm{rel}}$ converges similar to the running coupling constant. For $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ the $1+1$ spectrum is immediately converged because the full energy spectrum is used for its construction.
@@
where $E_{V}(n)$ is the energy calculated with the potential $V$ in the Hilbert space determined by the one-body cutoff $n$; $C=g^{2}\braket{\Psi}{\sum\limits_{i<j}\delta(x_{i}-x_{j})}{\Psi}$ is the contact parameter [57]. As we have argued above, the running coupling constant approach leads to a fast convergence, and hence $E_{g(n)}(n)$ can be considered constant with $n$ for sufficiently large cutoffs. We conclude that $E_{g}(n)$ features a slow $1/\sqrt{n}$ convergence for any number of particles, in agreement with the numerical analysis of Ref. [13].
@@
$E(n)=E(n\to\infty)+\frac{A}{n^{\sigma}}.$ (25)
The fit shows that renormalized interactions lead to a faster convergence rate (larger $\sigma$). It allows us to conclude that (i) the convergence of the energy is significantly faster for renormalized interactions in agreement with previous studies [34, 37, 20]; (ii) the effective interaction $V_{\mathrm{eff}}^{\mathrm{rel}}$ performs similar to the running coupling constant approach; (iii) the effective potential $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ outperforms other approaches.
Finally, we note that some energies for the $1+3$ system in Fig. 3 can converge from below, i.e., the ground-state energy increases when we increase the cutoff parameter. This is at odds with the behavior expected for CI methods, which are variational, but it is in fact a typical feature of renormalized interactions. To rationalize this, let us consider the running coupling constant, see Fig. 2. When the Hilbert space is extended, the attractive interaction effectively weakens, see Eq. (8). This can result in a competition between a decrease in energy due to a larger $n$ and an increase in energy due to a weaker potential [^9] Note that we cannot rule out an oscillatory convergence behavior, even though it was not observed in our data. In general, we expect potential oscillations to be suppressed for large cutoffs.. The non-variational character of our calculations is not necessarily a disadvantage. Instead, this can be used to approximate the error imposed by the finite basis: Deviations between the different interactions, in particular if one converges from above and one from below, allow for an estimate of the ‘true’ value that would result from an infinite Hilbert space.

### v2

@@ I Introduction
Experimental advances in cold-atom physics allowed one to create and study (quasi) one-dimensional (1D) systems of a few (up to 10) particles in a highly controllable setting [1, 2, 3]. This rapid progress fueled studies of the corresponding Schrödinger equation [4, 5]. Thanks to the small number of particles, this equation can be solved accurately using numerical methods, such as the multi-layer multi-configuration time-dependent Hartree method for atomic mixtures [6, 7, 8], flow equation approaches [9, 10, 11], coupled cluster methods [12, 13, 14], or exact diagonalization methods, also known as the configuration interaction method (CI) [15, 16, 17, 18, 19, 20, 21, 22, 23] and its extension, the importance truncated CI [24, 25]. These powerful numerical techniques, however, usually face problems for strong interactions (see, e.g., [17, 13]).
@@
Zero-range interactions require no regularization in one-dimensional geometries (unlike in 2D and 3D), however, Refs. [30, 31, 32, 33, 34, 35, 36, 37] demonstrate that renormalized interactions significantly boost the convergence of numerical simulations, enabling calculations with larger particle numbers and stronger interactions. However, to the best of our knowledge, these potentials were only benchmarked against the bare interaction. A direct comparison between different renormalization methods is still lacking and this paper aims to fill this gap. Additionally, we discuss a unitary transformation of observables together with the potential. This procedure is well-known in nuclear physics, but not in one-dimensional few-body problems. We will showcase that such a transformation can be necessary to improve the convergence of observables other than the energy.
@@
Our testbed for illustrating differences between different renormalization methods is a system of harmonically trapped fermions interacting via short range interactions. This system is well-studied (see, e.g., [15, 38, 16, 17, 13, 39]) and features slow convergence of observables [13], making it an ideal testbed for renormalized interactions and an optimal candidate for comparative studies. Furthermore, the corresponding two-body problem is analytically solvable [40, 41], which will facilitate construction of effective interactions in our calculations.
@@ II.1 System
where $x_{i}$ and $x_{j}$ are the coordinates of particles $i$ and $j$ in the laboratory frame, and $V(x_{i},x_{j})$ is their interaction potential. The operators of the kinetic energy $T$ and the external trapping potential $W$ are given by
@@
where $N=N_{\uparrow}+N_{\downarrow}$ is the number of fermions. We assume that the spin of a particle is conserved so that $N_{\uparrow}$ and $N_{\downarrow}$ are fixed. As we are interested in low-energy scattering of atoms, we assume that $V$ is a zero-range (contact) potential [46, 47]. This implies that only fermions of different spins can interact as the Pauli exclusion principle forbids two fermions with equal spin to occupy the same position.
@@ II.2 Numerical method: Configuration interaction method
where $H_{l}$ is the $l$th Hermite polynomial. In the CI method, we first construct antisymmetric $N-$particle basis states and then use them to build the Hamiltonian matrix, which is subsequently diagonalized using the Arnoldi/Lanczos algorithm [51].
@@ III Renormalized interaction
Three renormalizations of the contact interaction, i)-iii) are mentioned above in Sec. II.1 (see also Fig. 1). We will benchmark the convergence of the different regularization methods against the bare contact interaction and then compare them with each other.
@@ III.1 Running coupling constant
where $E_{1+1}^{\mathrm{unconf.}}$ is the two-body energy of the unconfined system. For now, we discuss only $g<0$ because we want to use the energy of a bound system with $E_{1+1}^{\mathrm{unconf.}}<0$ as the guiding condition for renormalization. Note that in 1D, there is always a bound state if $g<0$ [55].
By solving Eq. (6) in momentum space, we arrive at the following equation for the energy (see Sec. IV the supplementary material [56] for details):
@@
This equation renormalizes the interaction in a finite Hilbert space, in the sense, that the ground state energy $E_{1+1}^{\mathrm{unconf.}}$ becomes independent of the size of the Hilbert space. Instead of $g$, the two-body energy $E_{1+1}^{\mathrm{unconf.}}$ now determines the physical properties of the system.
In the limit of large $\lambda$, we derive from Eq. (8)
@@
where $C=g^{2}\braket{\Psi}{\delta(x)}{\Psi}$ is the contact parameter [57]. The functional behavior of Eq. (10) is in agreement with Refs. [58, 36]. We will see that, in general, this behavior depends on the trapping potential. In an unconfined system the cutoff $\lambda$ is a continuous parameter while for the trapped system the cutoff $n$ is discrete. For large quantum numbers [^3] Large means that the energy of an excited state is much larger than the excitation energy to a neighboring state of a higher energy. For the harmonic oscillator, which we discuss below, this means that $n\gg 1$. A different line of argument with the same conclusion can be based on the asymptotic approximation of Hermite polynomials, see, e.g., Ref. [59]., we connect $\lambda$ and $n$ below using the semi-classical approximation.
Figure 2: Running coupling constant for $g(n\to\infty)=-2.5067$ (corresponding to $E_{1+1}=-1$, the two-body energy in the trapped system). Dots show $g(n)$ calculated by explicitly fixing the two-body ground state energy of two fermions in a harmonic trap for each single-body cutoff $n$. Squares show $g_{\mathrm{unconf}.}$ calculated by integrating out high momentum contributions of two fermions in free space, Eq. (8). The parameter $\lambda$ is related to the single-body cutoff $n$ via Eq. (12). The solid line represents the value of the coupling constant for an infinitely large Hilbert space, $g(n\to\infty)=-2.5067$.
@@
For a derivation of this equation and an extension to general polynomial potentials, see Sec. V in the supplementary material [56]. We see that convergence is relatively slow (because $1/\lambda=1/(2\sqrt{n})$) for a harmonic potential (see also Ref. [13]), making the use of renormalized interactions for trapped systems more important than in an unconfined case.
To compare $g(n)$ with $g_{\mathrm{unconf.}}(2\sqrt{n})$, we fix the value of $g(n\to\infty)$ and insert the expression $E_{1+1}^{\mathrm{unconf.}}=-g(n\to\infty)^{2}/4$ in Eq. (8) to calculate $g_{\mathrm{unconf.}}$. Then, we compute the two-body ground state energy of the trapped system (see Eq. (23)) and tune $g(n)$ such that this energy is reproduced in a truncated Hilbert space for each cutoff. In Fig. 2 we show the calculated running coupling constants $g(n)$ and $g_{\mathrm{unconf.}}(2\sqrt{n})$ [$g(n\to\infty)=-2.5067$ corresponding to $E_{1+1}=-1$, the two-body energy in the trapped system]. The behavior of $g(n)$ resembles that of $g_{\mathrm{unconf.}}$ for all values of $n$ qualitatively. For large values of $n$ they agree even quantitatively (less than 1% disagreement for $n\geq 21$). The disagreement for small values of $n$ can be explained by the validity of the semi-classical approximation used to derive Eq. (12), as it only holds in the limit of large quantum numbers.
@@ III.2 Effective interactions
with $T$ the kinetic energy operator. This renormalization scheme could be interpreted as a unitary transformation of the Hamiltonian [29, 60], see also Sec. III in the supplementary material [56]. Therefore, operators should be transformed alongside the Hamiltonian transformation.
@@
Note that $\widetilde{O}^{(2)}_{\mathcal{M}\mathcal{M}}$ is represented in eigenstates of the two-body solution |$\ket{\Psi_{i}}$ while $O^{(2)}_{\mathcal{M}\mathcal{M}}$ is written in the two-body basis functions |$\ket{i}$ (cf. Sec. III B. in the supplementary material [56]).
@@
Since we know the exact two-body solution of our system in terms of relative and center-of-mass coordinates (see below) the transformation of observables is straightforward if an observable can be factorized in these coordinates. However, if operators do not factorize, the transformation requires more involved calculations, for details see Secs. III C. and D. in the supplementary material [56].
@@
with $x=(x_{1}-x_{2})/\sqrt{2}$ and $X=(x_{1}+x_{2})/\sqrt{2}$ being the relative and center-of-mass coordinates. Only the relative part of the wavefunction is influenced by the interaction and therefore $\Phi_{n_{i}}(X)$ is simply the harmonic oscillator eigenfunction; the corresponding energy is $E_{n_{i}}=n_{i}+1/2$. For the eigenfunction in relative coordinates, odd states are not influenced by the interaction; they are the ones of the harmonic oscillator with the same energy. For even states one derives:
@@
We use the general framework and the two-body solution presented above to build the effective interaction in two different ways. First, we construct the effective interaction in relative coordinates (i.e., using $\psi_{\nu_{i}}$ from above) and transform this interaction matrix to the laboratory frame (the frame in which Eq. (1) is defined) where our calculations are performed. This is possible as only the relative part of the wavefunction is affected by the interaction. This approach is the one used in Refs. [20, 30, 31]. For the relevant equations we refer to Sec. I in the supplementary material [56] and Ref. [20].
Note that $V_{\mathrm{eff}}^{\mathrm{rel}}$ does not lead immediately (i.e., for any cutoff parameter) to a correct full two-body spectrum in a truncated Hilbert space. Indeed, the Hamiltonian matrix built upon the truncated basis allows for coupling between relative and center-of-mass coordinates. As a result, the two-body spectrum is not exact when we construct the effective interaction in relative coordinates and then transform it to the laboratory frame. It is worth noting that for the low-energy part of the spectrum, the coupling between the relative and center-of-mass coordinates is in general weak; it decreases as the cutoff parameter increases.
@@ III.2.2 Effective interaction from the full two-body energy spectrum: $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$
We also build the effective interaction using the full two-body solution presented in Eq. (21), see the relevant equations in Sec. II of the supplementary material [56]. This interaction explicitly accounts for the coupling between the relative and center-of-mass coordinates in a truncated Hilbert space, and fixes the spectrum of the $1+1$ system to the exact values for all one-body cutoffs. Note that $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ is harder to construct in comparison to $V_{\mathrm{eff}}^{\mathrm{rel}}$, in particular, for $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ the overlap between the laboratory frame and the exact two-body solution has to be calculated. However, its performance is in general better compared against our other renormalization methods as we illustrate in the next section. Furthermore, $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ is the only option for non-harmonic traps for which a factorization into relative and center-of-mass motions is not possible.
Note that effective interactions discussed above can be constructed in a similar manner for higher-dimensional systems and other (non-harmonic) traps  [^7] When working with non-harmonic traps, one can employ harmonic oscillator eigenfunctions as one-body basis. In this approach, the effective interactions become independent of the trap and are constructed exactly as described in this paper. The information about the trap is then hidden in the one-body part of the Hamiltonian $A_{ij}$, which will be off-diagonal.. The prerequisite is accurate knowledge of a solution to a two-body problem, which can be obtained either analytically or numerically. A numerical solution is typically the only option for non-contact interactions and non-harmonic traps for which an analytical solution is not attainable.
@@ IV Results
Note that although we use the $1+1$ system to renormalize our interactions, not all two-body observables, for example the density, are fixed by the renormalization scheme, i.e., a large number of basis states might still be needed to obtain accurate results. Therefore, already the simplest system with one spin-up and one spin-down fermion provides valuable insight into the convergence properties of the different renormalization schemes. In particular, since the $1+1$ system is exactly solvable, we can use its exact solution for accurate benchmarking. In practical calculations, the simulation of larger particle numbers is more important. We showcase the convergence of the renormalized interactions by studying a $1+3$ system with $N_{\uparrow}=1$ and $N_{\downarrow}=3$. [We observed similar convergence for other particle numbers, in particular for larger systems, see App. D.]
@@ IV.1 Energy
Figure 3: Energy of the system for different particle numbers and interactions. Panels a) and b) demonstrate the energy of the second excited state of the $1+1$ system; panels c) and d) show the ground state energy of $1+3$. Panels a) and c) are for attractive interaction with $g=-2.5067$ (which corresponds to $E_{1+1}=-1$); b) and d) illustrate the system with repulsive interactions, $g=3$ (which corresponds to $E_{1+1}=1.6$). Numerical data are shown with markers of different shape; lines are added to guide the eye. The interested reader can see these data plotted as a function of $1/n$ and also using the log-log scale for $1+1$ in the supplementary material [56], Figs. S1 and S6.
The energies, $E(g)-E(g=0)$, obtained with the different renormalization schemes are shown in Figure 3 as a function of the one-body cutoff parameter, $n$. As a relevant scale to judge the convergence we use the harmonic oscillator energy unit, $\hbar\omega=1$; all data are shown with respect to it. Panels a) and b) present results for the second excited state [^9] We do not show the ground state energy of the $1+1$ system as this quantity is fixed for $g(n)\delta$ and $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$. of the $1+1$ system, for representative attractive and repulsive interactions, respectively. Panels c) and d) demonstrate the ground-state energy of the $1+3$ system, which does not have an analytical solution. Hence, its numerical solution is more relevant from the practical point of view.
We compare our numerical data for the $1+1$ system with the exact values, see Eq. (23). First thing to notice is that the results obtained with the bare interaction always converge slower to the limit $n\to\infty$ than those for the effective interactions. For example, for the attractive interaction the energy calculated with $g\delta(x)$ has $E(n=31)-E(n=\infty)\approx 0.33$ for the highest presented cutoff; at the same time, the energy calculated with $g(n)\delta(x)$ deviates from the analytical result by $E(n=31)-E(n=\infty)\approx 0.035$. For the effective interactions we see that $V_{\mathrm{eff}}^{\mathrm{rel}}$ converges similar to the running coupling constant. For $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ the $1+1$ spectrum is immediately converged because the full energy spectrum is used for its construction.
@@
where $E_{V}(n)$ is the energy calculated with the potential $V$ in the Hilbert space determined by the one-body cutoff $n$; $C=g^{2}\braket{\Psi}{\sum\limits_{i<j}\delta(x_{i}-x_{j})}{\Psi}$ is the contact parameter [57]. As we have argued above, the running coupling constant approach leads to a fast convergence, and hence $E_{g(n)\delta}(n)$ can be considered constant with $n$ for sufficiently large cutoffs. We conclude that $E_{g\delta}(n)$ features a slow $1/\sqrt{n}$ convergence for any number of particles, in agreement with the numerical analysis of Ref. [13].
@@
$E_{V}(n)=E_{V}(n\to\infty)+\frac{A_{V}}{n^{\sigma_{V}}}.$ (25)
Note that the parameters $E_{V}(n\to\infty),A_{V}$ and $\sigma_{V}$ depend on the potential $V$. For better readability, we will omit the subscript $V$ in the following and instead state explicitly to which potential these parameters belong. The fit shows that renormalized interactions lead to a faster convergence rate (larger $\sigma$). It allows us to conclude that (i) the convergence of the energy is significantly faster for renormalized interactions in agreement with previous studies [34, 37, 20]; (ii) the effective interaction $V_{\mathrm{eff}}^{\mathrm{rel}}$ performs similar to the running coupling constant approach; (iii) the effective potential $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ outperforms other approaches.
Finally, we note that some energies for the $1+3$ system in Fig. 3 can converge from below, i.e., the ground-state energy increases when we increase the cutoff parameter. This is at odds with the behavior expected for CI methods, which are variational, but it is in fact a typical feature of renormalized interactions (see also a non-monotonic behavior of the data calculated with $V_{\mathrm{eff}}^{\mathrm{rel+cm}}$ in Figs. 567). To rationalize this, let us consider the running coupling constant, see Fig. 2. When the Hilbert space is extended, the attractive interaction effectively weakens, see Eq. (8). This can result in a competition between a decrease in energy due to a larger $n$ and an increase in energy due to a weaker potential [^10] Note that we cannot rule out an oscillatory convergence behavior, even though it was not observed in our data. In general, we expect potential oscillations to be suppressed for large cutoffs.. The non-variational character of our calculations is not necessarily a disadvantage. Instead, this can be used to approximate the error imposed by the finite basis: Deviations between the different interactions, in particular if one converges from above and one from below, allow for an estimate of the ‘true’ value that would result from an infinite Hilbert space.